# Rajiv Raman

### v2

@@ 4 Contributions
We give polynomial time algorithms to construct an outerplanar intersection support for an outerplanar non-piercing system $(G,\mathcal{H},\mathcal{K})$. For a non-piercing graph system of treewidth $t$, we show that the supports for the primal and dual hypergraphs have treewidth $O(2^{t})$ and $O(2^{4t})$ respectively, and the intersection support has treewidth $2^{O(2^{t})}$. For the primal and dual settings, we show that the exponential blow-up in the treewidth of the support is sometimes necessary.
Theorem 6.
This also implies the existence of primal and dual supports for an outerplanar non-piercing graph system.
@@
The rest of the paper is organized as follows. In Section 5, we present our results on outerplanar graphs. In Section 6, we present our results on bounded treewidth graphs. We conclude in Section 8.
@@ 5 Outerplanar Support
In this section, we give a polynomial time algorithm that computes an intersection support for an outerplanar non-piercing intersection system.
@@ 5.1 Cycle $axax$-free systems
For a cycle $C$ embedded in the plane, and collections $\mathcal{H}$ and $\mathcal{K}$ of subgraphs (not necessarily connected) of $C$, we call $(C,\mathcal{H})$ a cycle system and $(C,\mathcal{H},\mathcal{K})$ a cycle intersection system. We start with the notions of cycle systems that (strong) $axax$-free and $abab$-free.
@@
Let $(C,\mathcal{H})$ be a cycle system. $H,H^{\prime}\in\mathcal{H}$ are an $axax$-pair if there are four distinct vertices $a_{1},x_{1},a_{2},x_{2}$ in cyclic order on $C$ such that $a_{1},a_{2}\in H\setminus H^{\prime}$ and $x_{1},x_{2}\in H^{\prime}$. $(C,\mathcal{H})$ is $axax$-free if there are no $axax$-pairs in $\mathcal{H}$.
Definition 11 ($abab$-free).
Clearly, if $(C,\mathcal{H})$ is $axax$-free, then it is $abab$-free. We need one more definition, that of strong $axax$-free cycle systems.
@@
  1. 1.
    For two families $\mathcal{H},\mathcal{K}$ of subgraphs of $C$, the intersection system $(C,\mathcal{H},\mathcal{K})$ satisfies the $axax$-free property if both $(C,\mathcal{H})$ and $(C,\mathcal{K})$ are $axax$-free.
  2. 2.
    $(C,\mathcal{H},\mathcal{K})$ is said to satisfy the intersection property if for any $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with four vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order on $C$ such that $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$, then $H\cap K\neq\emptyset$.
  3. 3.
    If $(C,\mathcal{H},\mathcal{K})$ satisfies the $axax$-free property and the intersection property, then it is said to satisfy the strong $axax$-free property.
@@
To see that $(C,\mathcal{H},\mathcal{K})$ satisfies the intersection property, consider $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order s.t. $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$. Since $H$ is connected in $G$, there is a path $P$ between $h_{1}$ and $h_{2}$, all of whose vertices lie in $H$. If $H\cap K=\emptyset$, then $P$ lies in $H\setminus K$. Since $h_{1}$ and $h_{2}$ are non-consecutive on $C$, $G\setminus P$ gets separated into two components with $k_{1}$ and $k_{2}$ in distinct components. This implies there is no path between $k_{1}$ and $k_{2}$ in $G$ that lies entirely in $K$, a contradiction since $K$ is connected.
Figure 1: Non-piercing implies $axax$-free. The figure is from [raman2025supportsgraphsboundedgenus].
@@
Let $C$ be the outer cycle in an outerplanar embedding of $G$. By Lemma 13, $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. By Theorem 14, there is an outerplanar support $\tilde{Q}$ for $(C,\mathcal{H},\mathcal{K})$. $\tilde{Q}$ is also a support for $(G,\mathcal{H},\mathcal{K})$ since the underlying intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ remains the same on $G$ and on $C$.
@@
If $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free, removing an $\mathcal{H}$-vertex or a $\mathcal{K}$-vertex leaves the reduced system $(C^{\prime},\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ is strong $axax$-free and the resulting intersection hypergraph remains the same since we did not remove any vertex $v\in H\cap K$ for any $H\in\mathcal{H}$ and $K\in\mathcal{K}$. Hence, a support for the reduced system is also a support for the original system.

### v3

@@ 4 Contributions
We give polynomial time algorithms to construct primal, dual, and intersection supports for hypergraphs defined by non-piercing subgraphs of a host graph that is outerplanar or has bounded treewidth. For a non-piercing graph system of treewidth $t$, we show that the supports for the primal and dual hypergraphs have treewidth $O(2^{t})$ and $O(2^{4t})$ respectively, and the intersection support has treewidth $2^{O(2^{t})}$.
@@
For the primal and dual settings, we show that the exponential blow-up in the treewidth of the support is sometimes necessary.
Theorem 10.
For any $m\in\mathbb{N}$, there is a non-piercing graph system $(G,\mathcal{H})$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ s.t. $\textsc{tw}(G)\leq m$, but $\textsc{tw}(Q)\geq\frac{2^{\lfloor m/2\rfloor}}{\sqrt{m}}$ for any primal support $Q$.
Theorem 11.
For any $m\in\mathbb{N}$, there is a non-piercing graph system $(G,\mathcal{H})$ s.t. $\textsc{tw}(G)\leq m$ and for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq\frac{2^{\lfloor m/2\rfloor}}{\sqrt{m}}$.
In case the host graph $G$ is outerplanar, we show that the non-piercing intersection system $(G,\mathcal{H},\mathcal{K})$ admits an outerplanar support.
Theorem 12.
Theorem 12 also implies the existence of primal and dual supports for an outerplanar non-piercing graph system. The authors in [buchin2011planar] left it open, the problem of deciding if a hypergraph admits an outerplanar support. While we don’t answer their question, we believe that the tools we develop for constructing outerplanar support can eventually help in resolving this decision problem
@@
The rest of the paper is organized as follows. In Section 5, we present the construction of support for bounded treewidth graphs systems, and in Section 6, we show lower bounds for the treewidth of primal and dual supports. In Section 7, we present results on outerplanar graphs. We conclude in Section 8.
5 Support of Bounded Treewidth
Lemma 13.
Proof.
The proofs follow immediately from the fact that $\mathcal{H}$ are non-piercing and the adhesion set corresponding to any edge in $E(T)$ is a separator in $G$. ∎
Lemma 14.
Proof.
Since $G$ is chordal, each bag induces a complete subgraph of $G$. Hence, each $H_{x}\in\mathcal{H}_{x}$ is connected. If $x$ is not a root of $T$, let $y$ be the parent of $x$. Then, for $H_{x},H^{\prime}_{x}\in\mathcal{H}_{x}$, the subgraph $H_{x}\setminus H^{\prime}_{x}$ of $G_{x}$ is connected since $H\setminus H^{\prime}$ is connected and $A_{xy}$ is a separator in $G$. Therefore, $(G_{x},\mathcal{H}_{x})$ is a non-piercing graph system. Finally, that $(T_{x},\mathcal{B}_{x})$ is a tree decomposition, is straightforward. ∎
5.1 Primal Support
Lemma 15.
Proof.
To show that $Q$ is a support, consider an $H\in\mathcal{H}$. Since $H$ is a connected subgraph in $G$, the vertices in $V(H)$ lie in bags corresponding to a connected sub-tree of $T$. Further, as $(T,\mathcal{B})$ is an easy tree decomposition, $\mathbf{b}(V(H))$ lie in bags of $\mathcal{B}^{\prime}$ corresponding to a connected sub-tree of $T$. Since each $B^{\prime}\in\mathcal{B}^{\prime}$ induces a complete subgraph of $Q$, it implies $V(H)$ induces a connected subgraph of $Q$. ∎
Lemma 16.
Let $(G,\mathcal{H})$ be a non-piercing graph system with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$. A tree decomposition $(T,\mathcal{B})$ of $G$ of width $t$ can be transformed into an easy tree decomposition $(T,\mathcal{B}^{\prime})$ of width at most $2^{t+2}+t$.
Proof.
We claim that $(T,\mathcal{B}^{\prime})$ is an easy tree decomposition. Suppose not. Let $H\in\mathcal{H}$ s.t. $H\cap B_{\rho}\neq\emptyset$ and $H\cap\mathbf{b}(B^{\prime}_{\rho})=\emptyset$. Then, $H\cap A_{x\rho}\neq\emptyset\neq H\cap A_{y\rho}$ by assumption. Let $H\cap A_{x\rho}=S$ and $H\cap A_{y\rho}=S^{\prime}$. Let $H^{\prime}$ be the minimal subgraph in $\mathcal{M}_{S}$ whose blue vertex was added to $B^{\prime}_{\rho}$ and let $H^{\prime\prime}$ be the minimal subgraph in $\mathcal{M}_{S^{\prime}}$ whose blue vertex was added to $B^{\prime}_{\rho}$. Then, $H$ and $H^{\prime}$ are incomparable in $(G_{x},\preceq)$. By part $(ii)$ of Lemma 13, this implies $H$ and $H^{\prime}$ are identical in $G_{-x}$. Similarly, $H$ and $H^{\prime\prime}$ are incomparable in $(G_{y},\preceq)$ and by Lemma 13 $H$ and $H^{\prime\prime}$ are identical in $G_{-y}$. It follows that $H^{\prime}|_{x}$ and $H^{\prime\prime}|_{x}$ are incomparable in $G_{x}$, and $H^{\prime}|_{y}$ and $H^{\prime\prime}|_{y}$ are incomparable in $G_{y}$. Also, note that $H^{\prime\prime}\cap A_{x\rho}=S=H^{\prime}\cap A_{x\rho}$. But this implies $H^{\prime}\setminus H^{\prime\prime}$ has at least two connected components - one in $G_{x}$ and one in $G_{y}$ contradicting the assumption that the subgraphs in $\mathcal{H}$ are non-piercing. Hence, $(T,\mathcal{B}^{\prime})$ is an easy tree decomposition. ∎
See 7
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$ where $T$ is a binary rooted tree. If $(T,\mathcal{B})$ is an easy tree decomposition, we are done by Lemma 15. Otherwise, we transform it into an easy tree decomposition of width at most $2^{t+2}+t$ by Lemma 16. We then obtain a primal support $Q$ of the same treewidth by Lemma 15. ∎
5.2 Dual Support
Lemma 17.
Proof.
For a vertex $v\in V$, consider $\mathcal{H}_{v}$, the set of all subgraphs containing $v$. Since $(T,\mathcal{B})$ is a tree decomposition, at each node $x\in V(T)$ s.t. $v\in B_{x}$, $\mathcal{H}_{v}\subseteq B^{\prime}_{x}$, where $B^{\prime}_{x}$ is the bag at node $x$ in $\mathcal{B}^{\prime}$, i.e., corresponding to $B_{x}$. Since the vertices in $B^{\prime}_{x}$ are pairwise adjacent, it implies $\mathcal{H}_{v}$ induces a connected subgraph of $Q^{*}$. Hence, $Q^{*}$ is a dual support. ∎
Lemma 18.
Proof.
$|\textrm{unique}(\mathcal{H}^{\prime}_{A_{xz}}|_{x})|<2^{t+1}$ (1)
We now argue that $Q^{*}$ is a dual support for $(G,\mathcal{H})$. Consider a vertex $v\in V(G)$. The algorithm above ensures that each subgraph is pushed at most once. Suppose $H^{\prime}\in\mathcal{H}_{v}$ was pushed by $H$ so that its modified copy $H^{\prime\prime}=H^{\prime}\setminus H$ does not cover $v$. Then, $H\in\mathcal{H}_{v}$. Let $e=\{a,b\}$ be the connecting edge between $H$ and $H^{\prime\prime}$ such that $a\in H$ and $b\in H^{\prime\prime}$. Since $(T,\mathcal{B})$ is a valid tree decomposition, there is a bag $B$ containing both $a$ and $b$. Let $H_{1}$ be the unique representative of $H^{\prime\prime}$ in $\textrm{unique}(\mathcal{H}^{\prime})$. Since Algorithm $k$-SDS puts a complete graph on the subgraphs intersecting $B$, it implies $H$ and $H_{1}$ are adjacent in $Q^{\prime}$. In $Q^{*}$, we made $H^{\prime}$ adjacent to $H_{1}$. Hence, $Q^{*}$ is a dual support for $(G,\mathcal{H})$. ∎
See 8
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. If $(T,\mathcal{B})$ is $2^{4(t+1)}$-sparse, then we are done. Otherwise, by Lemma 18, we obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, of treewidth at most $2^{4(t+1)}$. ∎
5.3 Intersection Support
Lemma 19.
Proof.
See 9
Proof.
Since the subgraphs in $\mathcal{H}$ are non-piercing, by Lemma 18, we can obtain a collection $\mathcal{H}^{\prime}$ s.t. $(T,\mathcal{B}^{\prime})$ is $2^{4(t^{\prime}+1)}$-sparse with respect to $\textrm{unique}(\mathcal{H}^{\prime})$. Note that $(T,\mathcal{B}^{\prime})$ remains $\mathcal{K}$-easy w.r.t. the intersection system $(G,\textrm{unique}(\mathcal{H}^{\prime}),\mathcal{K})$. Now, by Algorithm $k$-SDS we obtain a tree decomposition of a graph $Q^{*}$ that by Lemma 17 is a dual support for $(G,\textrm{unique}(\mathcal{H}^{\prime}))$ and $\textsc{tw}(Q^{*})\leq 2^{4(t^{\prime}+1)}=2^{{2^{t+4}}+4(t+1)}$. Since $Q^{*}$ is obtained by $k$-SDS, it also an intersection support for $(G,\textrm{unique}(\mathcal{H}^{\prime}),\mathcal{K})$. By Theorem 8, we can extend $Q^{*}$ to a dual support $\tilde{Q}$ for $(G,\mathcal{H})$ such that $\textsc{tw}(\tilde{Q})=\textsc{tw}(Q^{*})$. Then $\tilde{Q}$ is also an intersection support for $(G,\mathcal{H},\mathcal{K})$ since the tree decomposition $(T,\mathcal{B}^{\prime})$ is $\mathcal{K}$-easy. ∎
5.4 Implementation
In this section, we show that our algorithms for the construction of supports of bounded treewidth run in polynomial time if the treewidth of the host graph is bounded. We argue by showing that if $\mathcal{H}$ is a collection of non-piercing subgraphs of a bounded treewidth graph $G$, then the VC-dimension of the corresponding set system defined by $(V(G),\mathcal{H})$ is bounded, which together with Sauer-Shelah lemma (Chapter 10 in [Mat02LecDiscGeom]) implies that $|\mathcal{H}|$ is bounded by $\textrm{poly}(n)$ where $n=|V(G)|$.
Theorem 20.
Proof.
Next, we define a set of subgraphs in $\mathcal{H}$ that intersect with both $S_{1}$ and $S_{2}$. Such a set of subgraphs must exist since $S$ is shattered by element of $\mathcal{H}$. Let $J_{i}\in\mathcal{H}$ be a subgraph such that $J_{i}\cap S=H_{i}\cup H^{\prime}_{i}$ for $i=1,2,\ldots k$. Then, each $J_{i}$ contains exactly $2\lfloor\frac{t+2}{2}\rfloor$ vertices of $S$ and hence, they are pairwise incomparable in $S$. Also, each $J_{i}$ intersects the adhesion set $A_{e}$ as they are connected subgraphs. Since the subgraphs in $\mathcal{H}$ are non-piercing, by part $(ii)$ of Lemma 13, $J_{i}$ and $J_{\ell}$ should be incomparable in $A_{e}$ for $i\neq\ell$. But there are only $\binom{t+1}{\lfloor\frac{t+1}{2}\rfloor}<k$ subsets of $A_{e}$ that are pairwise incomparable. This contradicts the fact that the set $S$ can be shattered. Hence, the VC-dimension of $(V(G),\mathcal{H})$ is at most $33t+$ 3. By Sauer-Shelah lemma (see [Mat02LecDiscGeom], Chap. 10) therefore, $|\mathcal{H}|\leq O(n^{3t+3})$. ∎
Now, we are ready to show the running time of our algorithms.
Theorem 21.
Proof.
6 Lower Bounds
In this section, we show that there exist non-piercing graph systems for which the treewidth of any primal or dual support is exponential in the treewidth of the host graph.
See 10
Proof.
Let $n=\lfloor m/2\rfloor$ and let $N=\binom{n}{\lfloor n/2\rfloor}$. We construct an $N\times N$ grid $B$ of isolated vertices $b_{i,j}$, $i=1,\ldots,N$ and $j=1,\ldots,N$. This constitutes the set of vertices colored $\mathbf{b}$. Let $R=\{r_{1},\ldots,r_{n}\}$ and $C=\{c_{1},\ldots,c_{n}\}$ be two sets of $n$ isolated vertices each. The vertices in $R\cup C$ are colored $\mathbf{r}$. Let =$\mathcal{R}=\{R_{1},\ldots,R_{N}\}$ be the collection of $\binom{n}{\lfloor n/2\rfloor}$ subsets of $R$ of size $\lfloor n/2\rfloor$. Similarly, let $\mathcal{C}=\{C_{1},\ldots,C_{N}\}$ be the $\binom{n}{\lfloor n/2\rfloor}$ subsets of $C$ of size $\lfloor n/2\rfloor$. Let the host graph $G$ be the complete bipartite graph with bipartition $(R\cup C)$ and $B$.
For $i=1,\ldots,N$ and $j=1,\ldots,N-1$, let $H_{i;j,j+1}$ be the subgraph induced on the vertices $\{b_{i,j},b_{i,j+1}\}\cup R_{i}\cup C_{j+1}$ for $i=1,\ldots,N$ and $j=1,\ldots,N-1$. Similarly, for each $i=1,\ldots,N-1$ and $j=1,\ldots,N$, let $H_{i,i+1;j}$ be the subgraph induced on the vertices $\{b_{i,j},b_{i+1,j}\}\cup R_{i+1}\cup C_{j}$. See Figure 1(a). Let =$\mathcal{H}=\{H_{i;j,j+1}:i=1,\ldots,N,j=1,\ldots,N-1\}\cup\{H_{i,i+1;j}:i=1,\ldots,N-1,j=1,\ldots,N\}$.
By construction, each subgraph in $\mathcal{H}$ contains exactly two vertices colored $\mathbf{b}$ corresponding to vertices in the grid that are consecutive row-wise or column-wise, and one subset from each of the collections $\{R_{1},\ldots,R_{N}\}$ and $\{C_{1},\ldots,C_{N}\}$. Since $G$ is a complete bipartite graph, for any $H\in\mathcal{H}$, the two blue vertices in $H$ are adjacent to all red vertices in $H$. Hence, $H$ is a connected subgraph of $G$.
Consider two distinct subgraphs $H,H^{\prime}\in\mathcal{H}$. Since $H$ and $H^{\prime}$ differ in either their row index or column index, $H\setminus H^{\prime}$ and $H^{\prime}\setminus H$ each contains at least one blue vertex. Suppose $H$ and $H^{\prime}$ differ in their row index. By construction, $H$ and $H^{\prime}$ are adjacent to distinct subsets of size $\lfloor n/2\rfloor$ in $R$. Hence, both $(H\setminus H^{\prime})\cap R$ and $(H^{\prime}\setminus H)\cap R$ are non-empty. Similarly, if $H$ and $H^{\prime}$ differ in their column index, then $(H\setminus H^{\prime})\cap C$ and $(H^{\prime}\setminus H)\cap C$ are non-empty. Since $H\setminus H^{\prime}$ contains at least one blue vertex of $H$ and $G$ is a complete bipartite graph, each red vertex in $H\setminus H^{\prime}$ is connected to the blue vertices in $H\setminus H^{\prime}$. On the other hand, if both blue vertices of $H$ are present in $H\setminus H^{\prime}$, then they are adjacent via a red vertex in $H\setminus H^{\prime}$ since $H\setminus H^{\prime}\cap R$ or $H\setminus H^{\prime}\cap C$ is non-empty and $G$ is a complete bipartite graph. Thus, $H\setminus H^{\prime}$ is connected. A symmetric argument implies $H^{\prime}\setminus H$ is connected and hence $\mathcal{H}$ is a set of non-piercing subgraphs of $G$. Since $G$ is a complete bipartite graph, each blue vertex in $H\setminus H^{\prime}$ is adjacent to each red vertex in $H\setminus H^{\prime}$. Thus, $H\setminus H^{\prime}$ is connected. Similarly, $H^{\prime}\setminus H$ is connected and hence, $(G,\mathcal{H})$ is a non-piercing graph system.
Since $G$ is a bipartite graph with $2n$ red vertices, $G$ has treewidth at most $2n\leq m$. There is a subgraph in $\mathcal{H}$ corresponding to each pair of blue vertices that are consecutive along a row of the grid or along a column of the grid. Therefore, any primal support contains a grid of size $N\times N$ as a subgraph. Since $N=\binom{n}{\lfloor n/2\rfloor}\geq\frac{2^{n}}{\sqrt{2n}}$, it follows that the treewidth of any primal support is at least $\frac{2^{n}}{\sqrt{2n}}\geq\frac{2^{\lfloor m/2\rfloor}}{\sqrt{m}}$. ∎
((a)) Construction of primal hypergraph
((b)) Construction of dual hypergraph
Figure 1: Construction of exp. lower bounds on treewidth for primal and dual supports. (a) Subgraph $H_{2;2,3}$ consists of vertices $\{b_{2,2},b_{2,3}\}\cup R_{2}\cup C_{3}$. (b) Subgraph $H_{1,2}$ consists of vertices $\{b_{1,3},b_{2,3},b_{1,4},b_{2,4}\}\cup R_{1}\cup C_{2}$.
See 11
Proof.
The construction is similar to that of the primal lower bound. Let $n=\lfloor m/2\rfloor$ and let $N=\binom{n}{\lfloor n/2\rfloor}$. Let $B$ be a $(2N+1)\times(2N+1)$ grid of isolated vertices $b_{i,j}$ for $i,j\in\{1,\ldots 2N+1\}$. Let $R$ and $C$ be two sets of $n$ isolated vertices in each. The host graph is the complete bipartite graph $G$ with bipartition $(R\cup C)$ and $B$. Let =$\mathcal{R}=\{R_{1},\ldots,R_{N}\}$ denote the $\binom{n}{\lfloor n/2\rfloor}$ subsets of $R$ of size $\lfloor n/2\rfloor$. $\mathcal{C}=\{C_{1},\ldots,C_{N}\}$ is defined analogously.
We define the subgraphs $H_{i,j}$ so that each subgraph contains 4 vertices of $B$ in a checkerboard pattern. More formally, the subgraphs $H_{i,j}$ are defined as follows. Let $i\in\{1,\ldots,2N\}$. If $i$ is odd, then for $j=1,\ldots,N$, $H_{i,j}$ is the induced graph on $b_{i,2j-1},b_{i+1,2j-1},b_{i,2j},b_{i+1,2j}$ and the subset $R_{i}$ of $R$ and $C_{j}$ of $C$. If $i$ is even, then for $j=2,\ldots,N$, $H_{i,j}$ is the induced subgraph of $G$ on the vertices $b_{i,2j},b_{i+1,2j},b_{i,2j+1},b_{i+1,2j+1}$ and the subsets $R_{i}$ and $C_{j}$. See Figure 1(b). For odd (even) $i$, the vertices of $H_{i,j}$ in $B$ are shown by red (blue) squares. Let $\mathcal{H}$ denote the set of subgraphs thus constructed.
Each subgraph $H_{i,j}\in\mathcal{H}$ is induced on four vertices of $B$ and a subset $R_{i}$ of $R$ and a subset $C_{j}$ of $C$. Since $G$ is a complete bipartite graph, it follows that each subgraph $H\in\mathcal{H}$ induces a connected subgraph of $G$. Consider any two subgraphs $H,H^{\prime}\in\mathcal{H}$. Since they differ in at least one of the row or column indices, it follows that $H\setminus H^{\prime}$ and $H^{\prime}\setminus H$ each contain at least three vertices of $B$. Further, if $H$ and $H^{\prime}$ differ in their row index, then they contain distinct subsets in $\mathcal{R}$. Similarly, if $H$ and $H^{\prime}$ differ in their column index, they contain distinct subsets in $\mathcal{C}$. Since no two subsets in $\mathcal{R}$ are contained in one another, it follows that $(H\setminus H^{\prime})\cap R$ and $(H^{\prime}\setminus H)\cap R$ are non-empty if $H$ and $H^{\prime}$ differ in their row index. Similarly, $(H\setminus H^{\prime})\cap C$ and $(H^{\prime}\setminus H)\cap C$ are non-empty if $H$ and $H^{\prime}$ differ in their column index. Since $G$ is a complete bipartite graph and $H\setminus H^{\prime}$ contain vertices of $B$ and vertices of either $R\cup C$, it follows that $H\setminus H^{\prime}$ is connected. Similarly, $H^{\prime}\setminus H$ is connected. Hence, $\mathcal{H}$ is a collection of non-piercing subgraphs of $G$.
Since $|R\cup C|=2n$, it follows that $\textsc{tw}(G)\leq 2n\leq m$. On the other hand, for $i=2,\ldots,2N$ and $j=2,\ldots,2N$, the vertices $b_{i,j}$ are contained in exactly two subgraphs in $\mathcal{H}$, and thus, they have to be adjacent in any dual support for $(G,\mathcal{H})$. This implies that any dual support $Q^{*}$ must contain a grid of size $N\times N$ as an induced subgraph and therefore, $\textsc{tw}(Q^{*})\geq N$. Since $N=\binom{n}{\lfloor n/2\rfloor}$, $\textsc{tw}(Q^{*})\geq\frac{2^{n}}{\sqrt{2n}}\geq\frac{2^{\lfloor m/2\rfloor}}{\sqrt{m}}$. ∎
@@ 7 Outerplanar Support
In this section, we give a polynomial time algorithm that computes an intersection support for an outerplanar non-piercing intersection system. Outerplanar graphs have treewidth at most 2. Let $G$ be an outerplanar graph and $\mathcal{H},\mathcal{K}$ be collections of non-piercing subgraphs of $G$. By Theorem 7 and Theorem 8, the graph system $(G,\mathcal{H})$ admits primal and dual supports of treewidth at most $22^{4}+$ 2 and $2^{12}$ respectively. Further, by Theorem 9, $(G,\mathcal{H},\mathcal{K})$ admits an intersection support of treewidth at most $2^{2^{7}}$. However, we show that in this restricted setting, i.e., when $G$ is an outerplanar graph, there is an intersection support that is outerplanar. This also implies outerplanar primal and dual supports.
@@ 7.1 Cycle $axax$-free systems
For a cycle $C$ embedded in the plane, and collections $\mathcal{H}$ and $\mathcal{K}$ of subgraphs (not necessarily connected) of $C$, we call $(C,\mathcal{H})$ a cycle system and $(C,\mathcal{H},\mathcal{K})$ a cycle intersection system. We start with the definitions of cycle systems that (strong) $axax$-free.
@@
Let $(C,\mathcal{H})$ be a cycle system. $H,H^{\prime}\in\mathcal{H}$ are an $axax$-pair if there are four distinct vertices $a_{1},x_{1},a_{2},x_{2}$ in cyclic order on $C$ such that $a_{1},a_{2}\in H\setminus H^{\prime}$ and $x_{1},x_{2}\in H^{\prime}$. $(C,\mathcal{H})$ is $axax$-free if there are no $axax$-pairs in $\mathcal{H}$. For two families $\mathcal{H},\mathcal{K}$ of subgraphs of $C$, the intersection system $(C,\mathcal{H},\mathcal{K})$ is $axax$-free if both $(C,\mathcal{H})$ and $(C,\mathcal{K})$ are $axax$-free.
Definition 23 (Intersection property).
The cycle intersection system $(C,\mathcal{H},\mathcal{K})$ is said to satisfy the intersection property if for any $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with four vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order on $C$ such that $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$, then $H\cap K\neq\emptyset$.
@@
The cycle intersection system $(C,\mathcal{H},\mathcal{K})$ is said to satisfy the strong $axax$-free property, if it is $axax$-free and it satisfies the intersection property.
@@
To see that $(C,\mathcal{H},\mathcal{K})$ satisfies the intersection property, consider $H\in\mathcal{H}$ and $K\in\mathcal{K}$ with vertices $h_{1},k_{1},h_{2},k_{2}$ in cyclic order s.t. $h_{1},h_{2}\in H$ and $k_{1},k_{2}\in K$. Since $H$ is connected in $G$, there is a path $P$ between $h_{1}$ and $h_{2}$, all of whose vertices lie in $H$. If $H\cap K=\emptyset$, then $P$ lies in $H\setminus K$. Since $h_{1}$ and $h_{2}$ are non-consecutive on $C$, $G\setminus P$ gets separated into two components with $k_{1}$ and $k_{2}$ in distinct components. This implies there is no path between $k_{1}$ and $k_{2}$ in $G$ that lies entirely in $K$, a contradiction since $K$ is connected. ∎
Figure 2: Non-piercing implies $axax$-free.
@@
Let $C$ be the outer cycle in an outerplanar embedding of $G$. By Lemma 25, $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. By Theorem 26, there is an outerplanar support $\tilde{Q}$ for $(C,\mathcal{H},\mathcal{K})$. $\tilde{Q}$ is also a support for $(G,\mathcal{H},\mathcal{K})$ since the underlying intersection hypergraph defined by $\mathcal{H}$ and $\mathcal{K}$ remains the same on $G$ and on $C$. ∎
@@
If $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free, removing an $\mathcal{H}$-vertex or a $\mathcal{K}$-vertex leaves the reduced system $(C^{\prime},\mathcal{H}^{\prime},\mathcal{K}^{\prime})$ is strong $axax$-free and the resulting intersection hypergraph remains the same since we did not remove any vertex $v\in H\cap K$ for any $H\in\mathcal{H}$ and $K\in\mathcal{K}$. Hence, a support for the reduced system is also a support for the original system. ∎